A: Assume that when adults with smartphones are randomly​ selected,48% use them in meetings or classes. If6 adult smartphone users are randomly​ selected, find the probability that exactl 2of them use their smartphones in meetings or classes

B) A pharmaceutical company receives large shipments of aspirin tablets. The acceptance sampling plan is to randomly select and test

36 ​tablets, then accept the whole batch if there is only one or none that​ doesn't meet the required specifications. If one shipment of

6000 aspirin tablets actually has 5​% rate of​ defects, what is the probability that this whole shipment will be​ accepted? Will almost all such shipments be​ accepted, or will many be​ rejected?

The probability that this whole shipment will be accepted is?

(Round to four decimal places as​ needed.
Answer #1

Ans:

A)Use binomial distribution with n=6 and p=0.48

P(x=2)=6C2*0.48^2*(1-0.48)^4=0.2527

B)Use binomial distribution with n=36 and p=0.05

P(accepted)=P(x<=1)

=P(x=0)+P(x=1)

=(1-0.05)^36+36C1*0.05*(1-0.05)^35

=0.4567

Only 45.67% will be accepted and many will be rejected.